OpenStudy (anonymous):

How do you get rid of the "i" in 1i, what about 13i?

OpenStudy (anonymous):

Who put the i in it? i^2=-1

OpenStudy (anonymous):

\[1i=\sqrt{-1}\] \[13i=13\sqrt{-1}\]

OpenStudy (anonymous):

you can use that radical sign..

OpenStudy (anonymous):

multply it by itself...? like 1i x 1i = i^2 = -1 like what chaguanas said...

OpenStudy (anonymous):

that could be.. you have many choices to get rid of i.